OpenStudy (anonymous):

If 5p = -100 + 3Q2, which of the following could Q possibly be?

OpenStudy (mathstudent55):

What is 3Q2? Is there an exponent?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

The symbol ^ means exponent. It would be easier to understand if you used it. Otherwise, we need to waste time trying to figure out what the problem is.

OpenStudy (anonymous):

\[q=\left(\begin{matrix}5p-100 \\3\end{matrix}\right)\] \[Q=\left(\begin{matrix}5p+100 \\ 3\end{matrix}\right)\] \[Q=\left(\begin{matrix}-5p+100 \\3 \end{matrix}\right)\] \[Q=\left(\begin{matrix}-5p-100 \\ 3\end{matrix}\right)\]

OpenStudy (anonymous):

ok sorry

OpenStudy (anonymous):

i know its not A

OpenStudy (mathstudent55):

Wait. this is not exponents. What is 3Q2?

OpenStudy (anonymous):

3Q^2?

OpenStudy (anonymous):

dont pay attention to the file i just sent

OpenStudy (mathstudent55):

\(5p = -100 + 3Q2\) \(5p +100 = 3Q2\) \(\dfrac{5p +100}{3} = Q2\) \(Q = \pm\sqrt{ \dfrac{5p +100}{3} }\)

OpenStudy (anonymous):

thank you